# A universal black-box quantum Monte Carlo approach to quantum phase transitions

### v2

@@ Appendix B Parity restricted PMR-QMC for the TFIM
In our transverse-field Ising model (TFIM) numerical experiments of Section III.2, we computed the ES and FS restricted to the positive parity subspace of the TFIM. We briefly discuss the details of our methodology. First, we study a rotated TFIM for which the transverse field term is $\sum_{i}Z_{i}$, Eq. 20. In doing so, the parity operator,
@@
commutes with the TFIM, $[H,\mathcal{P}]=0$. Hence, the parity $P=\pm 1$ is a good quantum number, and it is well known that the groundstate of the TFIM lies in the $P=+1$ parity sector by the Perron-Frobenius theorem [18].
For our rotated TFIM choice, $\mathcal{P}$ is purely diagonal. For both SSE-QMC [18] and PMR-QMC (see Eq. 23), computational basis elements |$\ket{z}$ are part of the QMC configuration. At any given time in simulation, then, one can query the eigenvalue of $\mathcal{P}\ket{z}$, which is the instantaneous parity $P_{\mathcal{C}}$. Note that, in fact, |$\ket{z}$ is a bitstring, so the parity can readily be obtained by determining if there are an odd number of $1$’s in |$\ket{z}$ or not. In the SSE-QMC approach [18], the authors simply keep track of $P_{\mathcal{C}}$ during simulation and post-select on measurement outcomes.
In contrast, we perform PMR-QMC sampling restricted to the positive parity subspace the entire simulation. This is possible because the set of PMR-QMC updates is standardized in terms of the permutation group that comprises $H$. A careful look at the updates for spin-1/2 systems [23] reveals that there are only two updates that can change a configurations parity: a classical update or a block swap. In our TFIM implementation [32], we simply choose an initial |$\ket{z}$ with positive parity and only allow classical updates or block swaps that preserve parity. Both can be altered to preserve parity without post-selection, which is therefore more efficient than the SSE approach.
In a standard classical update, a random bit is flipped. In our code, we simply flip two random bits instead, which preserves parity. In a standard block swap, the classical state is also updated. Specifically, the permutation string $S_{\bm{i}_{q}}$ is split into two $S_{\bm{i}_{q}}=S_{2}S_{1}$ for $S_{1}=P_{i_{k}}\cdots P_{i_{1}}$ and $S_{2}=P_{i_{q}}\cdots P_{i_{k+1}}$ for $k\in\{1,\ldots,q-1\}$ chosen randomly. The classical state at position $k$ is then |$\ket{z^{\prime}}=S_{1}\ket{z}$ with energy $E_{z^{\prime}}$ The updated block-swap configuration is then $\mathcal{C^{\prime}}=\{\ket{z^{\prime}},S_{1}S_{2}\}$, which swaps $S_{1}$ an $S_{2}$ while updating |$\ket{z}\rightarrow\ket{z^{\prime}}$. The parity |$\ket{z^{\prime}}$ can change based on $S_{1}$.
For the TFIM, however, each $P_{i_{l}}$ is just a 2-body $X$ string, i.e., $X_{1}X_{2}$, $S_{\bm{i}_{q}}=\mathds{1}$ if and only if each permutation appears an even number of times. Together, this means $S_{1}$ flips the parity of |$\ket{z^{\prime}}$ if and only if $k$ is odd. Hence, parity can be preserved by only sampling $k\in\{1,\ldots,q-1\}$ that is also even. This can easily be accomplished, and is what we do in our code [32]. Together, these small changes ensure we stay in the $P=+1$ parity sector during the entire PMR-QMC simulation.
Appendix C Additional details on generating special unitaries
@@
In addition, $U$ can easily be shown to be unitary whenever $||\bm{c}||_{2}=1$ and $\{Q_{k}\}_{k=1}^{L}$ all anti-commute (see Eq. (6) in [65]). The largest maximal set of anti-commuting Paulis is $12n+$ 1, and there is an $O(n)$ algorithm to generate a canonical set of this type [66]. Let $\mathcal{Q}$ denote this canonical set of $201$ Paulis for a 100–spin system, which, in particular, includes $X^{\otimes 100}$. We can now describe our algorithm to generate 10 random unitaries such that $H_{U_{i}}\equiv U_{i}H(\lambda)U_{i}^{\dagger}$ satisfies the four properties described in the main text.
@@
To our knowledge, there is no simple calculation to check for the sign problem for these models a priori. So in practice, we simply ran a PMR-QMC simulation for each model for $\beta=5$ and $\lambda=1,$ which approximately the critical point for this model. The PMR-QMC code in Ref. [42, 23] automatically tracks the average and variance of the sign of the QMC weight, so we then simply post-selected those models with the highest average sign. Empirically, we find the worst average sign in this set of 10 is $0.9983\pm 0.0006$ (this is $2\sigma$), so this subset is essentially sign problem free.
Appendix D Derivations of PMR-QMC estimators for pure diagonal or pure off-diagonal driving
We derive the ES and FS estimators for the $H_{1}\propto D_{0}$ case stated in Sections IV.2 and IV.3. Most notably, we prove Theorems 1 and 2. We then substantiate the claim that these same estimators also straightforwardly carry over to the $H_{1}\propto H-D_{0}$ case.
D.1 Derivation strategy
@@
for $w_{\mathcal{C}}=D_{(z,S_{\bm{i}_{q}})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ the generalized PMR-QMC Boltzmann weight. For convenience, we write $O_{\mathcal{C}}\mathrel{\widehat{=}}\langle O\rangle$ to be read “$O_{\mathcal{C}}$ estimates $\langle O\rangle$.” Our first goal is to derive an estimator for the ES when $H_{1}\propto D_{0}$.
@@
gives the ES (see Eq. 16) up to a multiplicative correction, i.e., $\chi_{E}^{H_{1}}=\chi_{E}^{D_{0}}/\lambda^{2}$ (or more generally just $\chi_{E}^{H_{1}}\propto\chi_{E}^{D_{0}})$. Thus, we focus on attention on estimating $\chi_{E}^{D_{0}}$.
Our strategy is to first derive estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)D_{0}\rangle$ in Section D.2. As we shall see, $E_{z}\mathrel{\widehat{=}}\langle D_{0}\rangle$ does not depend on $\tau$. Then, in Section D.3, we show $\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle$ can be evaluated exactly and analytically into a valid estimator via novel divided difference relations. Together, these results give an $O(q^{2})$ estimator for ES. In Section D.3, we further simplify our estimator into a constant time $O(1)$ ES estimator using additional novel divided difference relations. This completes most of the work of proving Theorem 1.
@@
gives the FS (see Eq. 17) up to a multiplicative correction, and our strategy remains similar. Namely, we show in Section D.5 that $\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ can also be evaluated exactly and analytically via novel divided difference relations. This results in an $O(q^{4})$ estimator for FS. By yet additional novel divided difference relations, we reduce this to $O(q^{3})$ operations in Section D.6. This l lcompletes most of the work of proving Theorem 2.
Finally, we show how all the estimators we derived for the $H\propto D_{0}$ case carry over straightforwardly to the $H_{1}\propto H-D_{0}$ case and with the same complexities in Section D.7 thus completing the justification for Theorems 1 and 2.
D.2 Estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)D_{0}\rangle$
[…]
@@ D.2 Estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)D_{0}\rangle$
The basic idea is simple: without the middle $D_{0}$ term, we can safely apply the Leibniz rule to the off-diagonal expansion since $e^{-\tau H}$ and $e^{-(\beta-\tau)H}$ commute. Accounting for the non-commuting $D_{0}$, however, is simple—we just pick up a factor of $E_{z_{j}}$ from the $D_{0}\ket{z_{j}}$ as shown. Hence, we find
@@
as claimed in Eq. 27.
D.3 Energy susceptibility in quadratic time
@@
Lemma 1 (Convolution lemma).
@@
$\mathcal{L}\left\{(f*g)(t)\right\}$ =$=\mathcal{L}\left\{f(t)\right\}\mathcal{L}\left\{g(t)\right\}$ (59)
@@
D.4 Energy susceptibility in constant time
@@
Proposition 1 (Re-scaling relation).
@@
Lemma 2 (A weighted, repeated argument sum simplification).
@@
Corollary 1 (Computing the weighted, repeated argument sum).
@@
D.5 Fidelity susceptibility in quartic time
@@
Proposition 3 (Repeated argument sum simplification).
@@
Lemma 3 is general for any multiset $[x_{0},\ldots,x_{q}]$ partitioned at arbitrary index $j$. As a specific special case of interest now, we consider $[x_{j+1},\ldots,x_{q}]$ replaced with $[x_{j},\ldots,x_{q}]$,
@@
where the last multiset now contains an explicit extra $x_{j}$ from the concatenation of $[x_{j},\ldots,x_{q},x_{m}]$ and $[x_{r},\ldots,x_{j}]$ which both contain an explicit $x_{j}$. ∎
@@
D.6 Fidelity susceptibility in cubic time
@@
Lemma 4 (A weighted, repeated argument, double, less-than sum simplification).
@@
Corollary 3.
@@
The estimator follows from applying Eq. (83) to the estimator form given in Eq. (80), taking special care to correctly translate the sum from $i,j=0$ to $q$ on arguments $[x_{0},\ldots,x_{q}]$ to the sum of $j,m=r$ to $q$ on arguments $[E_{z_{r}},\ldots,E_{z_{q}}]$. The estimate $O(q^{3})$ comes from the observation that, for each $r$, we must do $O(q^{2})$ work. ∎
D.7 Off-diagonal case follows from diagonal case
@@
Firstly, $\langle\Gamma\rangle=\langle H-D_{0}\rangle=\langle H\rangle-\langle D_{0}\rangle$ by linearity. From prior work (see Eq.63 in Ref. [26]), we know
@@
which has the full non-trivial $\tau$ dependence in the familiar diagonal correlator. As such, we can estimate $\langle\Gamma(\tau)\Gamma\rangle$ and ES and FS integrals thereof provided we can derive estimators for $\langle H^{2}\rangle$ and $\langle D_{0}H\rangle$. By off-diagonal expansion manipulations, we know (e.g., Eq. 65 of Ref. [26]),
@@
Each of the static estimators, i.e., for $\langle D_{0}\rangle,\langle H\rangle,\langle H^{2}\rangle,and\langle D_{0}H\rangle$, can be computed in constant time $O(1)$. Integrals of these $\tau$-independent quantities are also trivial. Hence, the entire complexity in evaluating $\langle\Gamma(\tau)\Gamma\rangle$ is $O(q^{2})$ inherited from the complexity of computing $\langle D_{0}(\tau)D_{0}\rangle$. Similarly, the complexity of evaluating ES and FS is thus $O(1)$ and $O(q^{3})$ via the estimators we derived in Appendix D. Namely, we have now finished our proof of Theorems 1 and 2 by showing that the non-trivial parts of the ES and FS estimators for both the $H_{1}\propto D_{0}$ and $H_{1}\propto H-D_{0}$ are the same and hence have the same complexity.
Appendix E Most general driving term estimators
@@
E.1 The PMR form of $H_{1}$
In the study of quantum phase transitions, we assume $H(\lambda)=H_{0}+\lambda H_{1}$ for some $[H_{0},H_{1}]\neq 0$. Regardless of the specifics, this Hamiltonian can always be in PMR form, $H(\lambda)=\sum_{j}D_{j}(\lambda)P_{j}$, where we choose $P_{0}=\mathbb{1}$ by convention. Here, we show that one can always write
@@ E.1 The PMR form of $H_{1}$
E.2 Estimating $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$
In the prior section, we showed that we can write $H_{1}=\sum_{j}\Lambda_{j}D_{j}P_{j}$ for $\Lambda_{j}$’s diagonal matrices whenever $H=\sum_{j}D_{j}P_{j}$ and $P_{0}=\mathds{1}$. By linearity, it is sufficient to derive estimators for each $\langle\Lambda_{j}D_{j}P_{j}\rangle$ term, for which it is useful to separate diagonal term, $\Lambda_{0}$ from the general off-diagonal terms $\Lambda_{l}D_{l}P_{l}$ for $l>0$. In both cases, some of the present authors have derived the estimators (see Eqs.56 and 82) [26],
$\Lambda_{0}(z)$ =^$\mathrel{\widehat{=}}\langle\Lambda_{0}\rangle$ (95)
$\delta_{P_{l}}^{(q)}\frac{\Lambda_{l}(z)e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}$ =^$\mathrel{\widehat{=}}\langle\Lambda_{l}D_{l}P_{l}\rangle,$ (96)
and hence, we can evaluate $\langle H_{1}\rangle$ in $O(R)$ calls to these $O(1)$ estimators. Here, $\Lambda_{l}(z)\equiv\langle z|\Lambda_{l}|z\rangle$ and $\delta_{P_{l}}^{(q)}=1$ if the $q^{\text{th}}$ permutation in $S_{\bm{i}_{q}}$ is $P_{l}$ and is 00 otherwise. (Recall that products of permutations $S_{\bm{i}_{q}}$ partially make up PMR-QMC configurations as explained in Appendix A.)

### v3

@@ Appendix C Parity restricted PMR-QMC for the TFIM
In our transverse-field Ising model (TFIM) numerical experiments of Section II, i.e. Fig. 2, we computed the ES and FS restricted to the positive parity subspace of the TFIM. We briefly discuss the details of our methodology. First, we study a rotated TFIM for which the transverse field term is $\sum_{i}Z_{i}$, Eq. (2). In doing so, the parity operator,
@@
commutes with the TFIM, $H,\mathcal{P}]=0$. Hence, the parity $P=\pm 1$ is a good quantum number, and it is well known that the groundstate of the TFIM lies in the $P=+1$ parity sector by the Perron-Frobenius theorem [4].
[…]
Appendix D Additional details on generating special unitaries
@@
In addition, $U$ can easily be shown to be unitary whenever $\bm{c}||_{2}=1$ and $\{Q_{k}\}_{k=1}^{L}$ all anti-commute (see Eq. (6) in [30]). The largest maximal set of anti-commuting Paulis is $2n+1$, and there is an $O(n)$ algorithm to generate a canonical set of this type [50]. Let $\mathcal{Q}$ denote this canonical set of $201$ Paulis for a 100–spin system, which, in particular, includes $X^{\otimes 100}$. We can now describe our algorithm to generate 10 random unitaries such that $H_{U_{i}}\equiv U_{i}H(\lambda)U_{i}^{\dagger}$ satisfies the four properties described in the main text.
@@
Figure 11: Here, we plot summary statistics of all 50 random models described generally above. Each plot shows how the average sign (for fixed QMC parameters) is correlated with important simulation properties like (left) simulation time (b) average value of $q$, and (c) percent error in estimating the FS. In Fig. 4, we only simulate those 10 random models with the highest average sign in this plot.
To our knowledge, there is no simple calculation to check for the sign problem for these models a priori. So in practice, we simply ran a PMR-QMC simulation for each model for $\beta=5$ and $\lambda=1,$ which approximately the critical point for this model. The PMR-QMC code in Ref. [45, 8] automatically tracks the average and variance of the sign of the QMC weight, so we then simply post-selected those models with the highest average sign. Empirically, we find the worst average sign in this set of 10 is $0.9983\pm 0.0006$ (this is $2\sigma$), so this subset is essentially sign problem free at $\lambda=1.0$ (but this does not hold for all $\lambda$ as seen before in Fig. 8). This post-selection step is contained in a Jupyter notebook of our code [20], namely in before_fig2_post_selection.ipynb. We provide a graphical summary of the practical meaning of this post selection process in Fig. 11. In short, the worse the sign problem is, the longer the simulation takes to run and the less accurate it is for a fixed set of QMC simulation parameters. By post-selecting on those models with the 10 highest average signs, we thus generate an ensemble with random, high weight Pauli terms that is nevertheless efficiently simulable in PMR-QMC.
Appendix E Derivations of PMR-QMC estimators for pure diagonal or pure off-diagonal driving
We derive the ES and FS estimators for the $H_{1}\propto D_{0}$ case stated in Section IV. Most notably, we prove Theorems 1 and 2. We then substantiate the claim that these same estimators also straightforwardly carry over to the $H_{1}\propto H-D_{0}$ case.
E.1 Derivation strategy
@@
for $w_{\mathcal{C}}=D_{(z,S_{\bm{i}_{q}})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ the generalized PMR-QMC Boltzmann weight. For convenience, we write $O_{\mathcal{C}}\mathrel{\widehat{=}}\langle O\rangle$ to be read $O_{\mathcal{C}}$ estimates $\langle O\rangle$.” Our first goal is to derive an estimator for the ES when $H_{1}\propto D_{0}$.
@@
gives the ES (see Eq. Eq. 20) up to a multiplicative correction, i.e., $\chi_{E}^{H_{1}}=\chi_{E}^{D_{0}}/\lambda^{2}$ (or more generally just $\chi_{E}^{H_{1}}\propto\chi_{E}^{D_{0}})$. Thus, we focus on attention on estimating $\chi_{E}^{D_{0}}$.
Our strategy is to first derive estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)D_{0}\rangle$ in Section E.2. As we shall see, $E_{z}\mathrel{\widehat{=}}\langle D_{0}\rangle$ does not depend on $\tau$. Then, in Section E.3, we show $\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle$ can be evaluated exactly and analytically into a valid estimator via novel divided difference relations. Together, these results give an $O(q^{2})$ estimator for ES. In Section E.3, we further simplify our estimator into a constant time $O(1)$ ES estimator using additional novel divided difference relations. This completes most of the work of proving Theorem 1.
@@
gives the FS (see Eq. (21)) up to a multiplicative correction, and our strategy remains similar. Namely, we show in Section E.5 that $\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ can also be evaluated exactly and analytically via novel divided difference relations. This results in an $O(q^{4})$ estimator for FS. By yet additional novel divided difference relations, we reduce this to $O(q^{3})$ operations in Section E.6. This completes most of the work of proving Theorem 2.
Finally, we show how all the estimators we derived for the $H\propto D_{0}$ case carry over straightforwardly to the $H_{1}\propto H-D_{0}$ case and with the same complexities in Section E.7 thus completing the justification for Theorems 1 and 2.
E.2 Estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)\,D_{0}\rangle$
Since $\langle z|D_{0}e^{-\beta H}|z\rangle=E_{z}\langle z|e^{-\beta H}|z\rangle$, then by a direct off-diagonal series expansion of Tr[$\Tr[D_{0}e^{-\beta H}]$ (see also Sec. VIII.A of Ref. [19]), we find
@@ E.2 Estimators for $\langle D_{0}\rangle$ and $\langle D_{0}(\tau)\,D_{0}\rangle$
[…]
@@
as claimed in Eq. (27).
E.3 Energy susceptibility in quadratic time
@@
Lemma 1 (Convolution lemma [66]).
@@
$\mathcal{L}\left\{(f*g)(t)\right\}$ $=\mathcal{L}\left\{f(t)\right\}\mathcal{L}\left\{g(t)\right\}$ (59)
@@
E.4 Energy susceptibility in constant time
@@
Proposition 1 (Re-scaling relation [66]).
@@
Lemma 2 (A weighted, repeated argument sum simplification [66]).
@@
Corollary 1 (Computing the weighted, repeated argument sum [66]).
@@
E.5 Fidelity susceptibility in quartic time
@@
Proposition 3 (Repeated argument sum simplification [66]).
@@
Lemma 3 is general for any multiset $x_{0},\ldots,x_{q}]$ partitioned at arbitrary index $j$. As a specific special case of interest now, we consider $x_{j+1},\ldots,x_{q}]$ replaced with $x_{j},\ldots,x_{q}]$,
@@
where the last multiset now contains an explicit extra $x_{j}$ from the concatenation of $x_{j},\ldots,x_{q},x_{m}]$ and $x_{r},\ldots,x_{j}]$ which both contain an explicit $x_{j}$. ∎
@@
E.6 Fidelity susceptibility in cubic time
@@
Lemma 4 (A weighted, repeated argument, double, less-than sum simplification [66]).
@@
Corollary 3 (Computing the complicated weighted, repeated arugment, double, less-than sum [66]).
@@
The estimator follows from applying Eq. (83) to the estimator form given in Eq. (80), taking special care to correctly translate the sum from $i,j=0$ to $q$ on arguments $x_{0},\ldots,x_{q}]$ to the sum of $j,m=r$ to $q$ on arguments $E_{z_{r}},\ldots,E_{z_{q}}]$. The estimate $O(q^{3})$ comes from the observation that, for each $r$, we must do $O(q^{2})$ work. ∎
E.7 Off-diagonal case follows from diagonal case
@@
Firstly, $\langle\Gamma\rangle=\langle H-D_{0}\rangle=\langle H\rangle-\langle D_{0}\rangle$ by linearity. From prior work (see Sec. VI.B in Ref. [19]), we know
@@
which has the full non-trivial $\tau$ dependence in the familiar diagonal correlator. As such, we can estimate $\langle\Gamma(\tau)\Gamma\rangle$ and ES and FS integrals thereof provided we can derive estimators for $\langle H^{2}\rangle$ and $\langle D_{0}H\rangle$. By off-diagonal expansion manipulations, we know (e.g., Sec. VI.B of Ref. [19]),
@@
Each of the static estimators, i.e., for $\langle D_{0}\rangle,\langle H\rangle,\langle H^{2}\rangle,and\langle D_{0}H\rangle$, can be computed in constant time $O(1)$. Integrals of these $\tau$-independent quantities are also trivial. Hence, the entire complexity in evaluating $\langle\Gamma(\tau)\Gamma\rangle$ is $O(q^{2})$ inherited from the complexity of computing $\langle D_{0}(\tau)D_{0}\rangle$. Similarly, the complexity of evaluating ES and FS is thus $O(1)$ and $O(q^{3})$ via the estimators we derived in Appendix E. Namely, we have now finished our proof of Theorems 1 and 2 by showing that the non-trivial parts of the ES and FS estimators for both the $H_{1}\propto D_{0}$ and $H_{1}\propto H-D_{0}$ are the same and hence have the same complexity.
Appendix F Most general driving term estimators
@@
F.1 The PMR form of $H_{1}$
In the study of quantum phase transitions, we assume $H(\lambda)=H_{0}+\lambda H_{1}$ for some $H_{0},H_{1}]\neq 0$. Regardless of the specifics, this Hamiltonian can always be in PMR form, $H(\lambda)=\sum_{j}D_{j}(\lambda)P_{j}$, where we choose $P_{0}=\mathbb{1}$ by convention. Here, we show that one can always write
@@ F.1 The PMR form of $H_{1}$
F.2 Estimating $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$
In the prior section, we showed that we can write $H_{1}=\sum_{j}\Lambda_{j}D_{j}P_{j}$ for $\Lambda_{j}$’s diagonal matrices whenever $H=\sum_{j}D_{j}P_{j}$ and $P_{0}=\mathds{1}$. By linearity, it is sufficient to derive estimators for each $\langle\Lambda_{j}D_{j}P_{j}\rangle$ term, for which it is useful to separate diagonal term, $\Lambda_{0}$ from the general off-diagonal terms $\Lambda_{l}D_{l}P_{l}$ for $l>0$. In both cases, some of the present authors have derived the estimators (see Table I in [19]),
$\Lambda_{0}(z)$ $\mathrel{\widehat{=}}\langle\Lambda_{0}\rangle$ (95)
and hence, we can evaluate $\langle H_{1}\rangle$ in $O(R)$ calls to these $O(1)$ estimators. Here, $\Lambda_{l}(z)\equiv\langle z|\Lambda_{l}|z\rangle$ and $\delta_{P_{l}}^{(q)}=1$ if the $q^{\text{th}}$ permutation in $S_{\bm{i}_{q}}$ is $P_{l}$ and is 0 otherwise. (Recall that products of permutations $S_{\bm{i}_{q}}$ partially make up PMR-QMC configurations as explained in Appendix A.)